Circular No. 6041

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

JUPITER AND COMET 1993e R. Prange and C. Emerich, Institut d’Astrophysique Spatiale, Orsay;

A. Talavera, International Ultraviolet Explorer (IUE) Observatory, European Space Agency, Vilspa; W. Harris, G. Ballester, and M. Combi, University of Michigan; T. Livengood, Goddard Space Flight Center, NASA; and M. McGrath, Space Telescope Science Institute, report: “Nearly continuous ultraviolet spectral observations of Jupiter (range 115–330 nm, resolutions 0.014–0.6 nm) have been performed with IUE during July 13–24. These had been preceded by extensive observations in June, which provided a comprehensive, two-dimensional baseline database (latitude/local time), to which the impact and post-impact observations can be compared. Using the 3′′ small aperture, we have monitored the evolution of the ultraviolet spectral features in the areas caused by impacts from comet fragments A, B, E, G, K, Q, R, S, and W as they transited across the Jovian disk. The observations have been repeated a few rotation periods later, and monitoring of impact sites will continue until Aug. 15. Many absorption and emission features have been seen. Preliminary comparison of spectra from impact sites with those taken under similar geometrical conditions before and after the impact of nucleus A shows new, unidentified absorption lines at 176 and 178 nm and emission lines near 182 and 330 nm. The first emission lines could be tentatively attributed to S I lines at 180.8/182.0 nm, or to a brightening of the pre-existing Si II triplet at 180.7/181.6/181.7 nm. The second one might be due to a Na fluorescence line at 330.2 nm. The comparison also suggests a global brightening of the impact region by 10– 20 percent. Although small pointing errors could simulate such an effect, it is consistent with the Hubble Space Telescope WFPC2 images reported for impact A. The brightness decreased to its normal level in < 2 hr, as the feature rotated across the disk. The auroral emission at Lyα and in the H_{2} Werner and Lyman bands has generally been weak (especially the northern aurora), possibly as a consequence of the passage of dust through the Jovian magnetosphere. The Fine Error Sensor was used to attempt detection of possible brightening of the Jovian satellites from the flash of the direct impacts; Europa and Io were thus observed for impacts A, H, L, and Q. No noticeable brightening at the 1-percent level was found for impacts A and H.”

1994 July 29 Daniel W. E. Green